1. The number of cinema Ticket in Karachi and the the ticket fares are regulated. The fare currently charged is Rs.500 a ticket. The Cinema Association want to obtain governmnet's permission to raise the fare to increase their reveniues and ask you to be their economic adviser. After Studying the market, you come up with the following demand schedule for cinema Tickets:
Price ( Rupees Per ticket) Quantity Demanded (ticket Per month)
300 160
400 120
500 80
600 40
700 0

(b) What happens to the cinema's total revenue if the fare rises from 500 to 600? Should the ticket sellers try to obtain permission to raise the fare.
1
Expert's answer
2020-12-15T11:24:06-0500

When

"P_1 = 500 \\

Q_1 = 80"

Total revenue

"TR_1= P_1 \times Q_1 \\

= 500 \times 80 \\

= 40000"

When

"P_2 = 600 \\

Q_2 = 40"

Total revenue

"TR_2= P_2 \times Q_2 \\

= 600 \times 40

= 24000"

Since, "TR_2 < TR_1" , the total earnings revenue declines due to rising in fare, so the seller should not try to obtain permission to raise the fare.